# Displacement correlations in a two-dimensional colloidal liquid and their relationship with shear strain correlations

### v1

@@ I Introduction
The formulations mentioned above are focused on the mobility of particles and not on their direction of motion. The importance of the directional aspect of motion was noticed by Doliwa and Heuer [11], who demonstrated that the displacement correlations in the longitudinal direction behave quite differently from those in the transverse direction. Later, it was noticed that the transverse displacement correlations contain information about glass elasticity [4, 12, 13]. Thus we are interested in tensorial correlations of displacements in glassy liquids, which provides a key topic of the present work.
@@
Motivated by the above considerations, here we concretize our interest in displacement-based statistical quantities by discussing two-particle displacement correlations (DC) and shear strain correlations (SC) in two-dimensional liquids. Our first question is whether and to what extent these correlations are detectable in liquids that are only slightly dense. We focus on such density that the lifetime of “bonds” between the nearest neighbors is barely non-zero and not very long. Although the liquid in such cases does not exhibits macroscopic elasticity at all, we will show numerically an appreciable presence of DC and SC at lengthscales greater than the particle diameter. The time dependence of these correlations will also be investigated. The second question concerns the possible equivalence between the DC and the SC, as both of them are calculable from the same data of particle displacements. Does the usage of the common ingredient mean that the two kinds of correlations are equivalent? We will answer this question negatively, showing that the SC can be calculated from the DC but not vice versa. In other words, the DCs are more informative than the SCs, which leads to the third question: what is the nature of the extra information contained in DC? This question will be answered, at least partially, by showing that the DCs behave logarithmically for shorter distances, with the coefficient of the logarithmic term being an increasing function of the density. We interpret the nature of this extra information as manifestation of the cage effect.
@@ II.2 Displacement in liquids
The cage effect suggests the possibility that, in dense liquids, hydrodynamic description may be extended to shorter lengthscales close to the interparticle distance [3, 23]. Hydrodynamics means finding a closed set of dynamical equations for slow variables. The long lifetime of the “bonds” can provide grounds for slow variables in dense liquids, such as the temporally coarse-grained velocity of a particle. We emphasize that the situation is different from that of low-density fluids, in which the hydrodynamic velocity field is introduced by spatial coarse-graining of momentum as a conserved quantity. In deriving the Navier–Stokes equation for Newtonian fluids, any confusion between the velocity of a fluid element and the molecular velocity should be severely criticized. In dense liquids, however, the cage effect makes it possible—probably as a crude approximation—to regard the velocity of a particle as the velocity of a “fluid element” comprising the particle, with the understanding that fluctuations at the collisional frequency should be filtered out mostly by time-averaging.
@@ II.3 Displacement correlations
Typical behavior of the DCs as functions of $\tilde{\mathbf{d}}$, with $t_{\Delta}$ fixed, is depicted in Fig. 1. Details of the particle model used here [20] will be specified later in Sec. IV. It is seen in Fig. 1(a) that the two components of DC behave quite differently: $X_{\parallel}$ is positive everywhere, while $X_{\perp}$ becomes negative for large $\tilde{{d}}$. This difference was noted in a pioneering work by Doliwa and Heuer [11], many years before the researchers started to notice that the transverse DC may contain information about shear modulus of glassy systems [13, 12]. The directional behavior of DCs can be illustrated pictorially as a pair of vortices [20, 11], as is exemplified in Fig. 1(b).
@@ II.4 Shear strain correlations
The four-petaled SC pattern is often interpreted as a sign of elasticity, showing the response of an elastic media to localized plastic events. Although the observations of the SC pattern in sheared glasses [17, 18] seems to be consistent with this interpretation, it is curious that non-vanishing SC is reported also in quiescent glassy liquids, being visible even for timescales longer than $\tau_{\alpha}$ [19].
@@ III Relations involving shear strain correlations
In order to answer this question, let us develop a theoretical framework to derive an analytical relation between the DC and the SC.
@@ III.2.3 Relation between DC and SC
$=-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{1^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{y}}(\bm{\xi}^{\prime})}\right\rangle$
$\qquad{}-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{2^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{x}}(\bm{\xi}^{\prime})}\right\rangle$ (38)
@@ III.3 Possible nonequivalence between DC and SC
With the relation in Eqs. (45c) we can calculate the SC from the DC, but this does not necessarily mean that SC and DC are equivalent. The problem is whether the relation is invertible, making it possible to obtain the DC from the SC. Intuitively speaking, since differentiation is involved in the operators $\hat{L}_{\pm}$, some information in DC is likely to be missing from SC.
The invertibility of the linear operators $\hat{L}_{\pm}$ in Eq. (45c) can be discussed by checking their null spaces. The null space of $\hat{L}_{+}$, also known as the kernel of the linear operator [35], is the set of all the solutions $h_{+}$ to the homogeneous equation
@@
with the domain of the function taken as some physically appropriate range of $\xi_{*}$. Noticing that $\hat{L}_{+}$ is essentially the radial part of the Laplacian operator in the $\bm{\xi}_{*}$-space, we can readily find $h_{+}=A+B\log\xi_{*}$ (with $A$ and $B$ denoting arbitrary constants); in other words, the null space is spanned by $\{1,\log\xi_{*}\}$. If $f_{+}\sim h_{+}$, it is mapped to zero by $\hat{L}_{+}$ in Eqs. (45c) and therefore the information of $f_{+}$ is lost from $\chi_{\gamma}^{(0)}$. This makes the DC and the SC nonequivalent.
@@
Thus we have derived analytical relations involving the SCs, which allows us to discuss the possibility that the DCs and the SCs are not equivalent due to the null space of $\hat{L}$. Now let us proceed to numerical study of these analytical results, beginning with specification of the particle system in the next section.
@@ IV Specification of the particle system
As the interaction potential $V_{jk}$ between the $j$-th and $k$-th particles (separated by the relative position vector $\mathbf{r}_{ij}$), we adopt the harmonic repulsive potential,
$V_{jk}=\begin{cases}V_{\text{max}}\left(1-\dfrac{\left|{\mathbf{r}_{ij}}\right|}{\sigma}\right)^{2}&(\left|{\mathbf{r}_{ij}}\right|<\sigma)\cr 0&(\text{otherwise})\end{cases}$
@@ V Numerical results
We performed numerical simulation of the system of Brownian disks specified in Sec. IV. The system contains $N=4000$ particles, and $L$ is adjusted so as to give three values of the area fraction: $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$. To mimic the hardcore interaction, we chose $V_{\text{max}}=50$ for $\phi_{\text{area}}=0.50$, $V_{\text{max}}=500$ for $\phi_{\text{area}}=0.60$, and $V_{\text{max}}=5000$ for $\phi_{\text{area}}=0.70$ in units of ${k_{\text{B}}}T$; the time step for numerical integration, $\Delta{t}$, is chosen so as to satisfy $(\Delta{t}/\tau_{0})({k_{\text{B}}}T/V_{\text{max}})=0.1$ where $\tau_{0}=\sigma^{2}/D$.
@@ VI.1 Elastic modeling
In parallel to the term with $D^{\text{c}}_{*}$ in Eq. (52a), the presence of the term with ${\mu_{\text{r}}}D^{\text{c}}_{*}$ in Eq. (52b) means introducing the shear modulus ${\mu_{\text{r}}}{k_{\text{B}}}T/S$ into the model. While this is a natural modeling of DCs in the case of glass solids with plateau modulus or with idealization of $\tau_{\alpha}\to\infty$ [10, 4, 12], it may seem questionable to what extent the elastic modeling is applicable to the case of glassy liquids with modest $\tau_{\alpha}$. Here we regard the elastic modeling as a convenient starting point for searching more sophisticated description of $C_{a}^{\Delta}$ with a wider range of applicability, involving solid-like elasticity at some scale. A recent experiment on confined liquid glyserol [39] may support relevance of such solid-like elasticity in liquids far away from glass transition.
Given the numerical data of the DCs, we focus our attention to the ratio of the coefficients $S$ and $S/{\mu_{\text{r}}}$ in Eq. (51). In principle, ${\mu_{\text{r}}}$ can be computed from the data in the Fourier space as the ratio of the saturation values of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$:
@@
$X_{\parallel}={\frac{1}{2}}\left({f_{+}}+{f_{-}}\right),\quad X_{\perp}={\frac{1}{2}}\left({f_{+}}+{f_{-}}\right)$ (56)
@@
The analytical expressions of $f_{\pm}$ in Eqs. (61) and (62) allow us to extract information of elasticity from the numerical data of the particle system, on the assumption that the elastic model is quantitatively valid. They may provide useful alternatives to Eq. (53), as $\mathsf{X}(\tilde{\mathbf{d}},t_{\Delta})$ is easier to compute than its Fourier counterpart used in Eq. (53).
@@
It is easy to verify that the above expressions for DCs and SCs are consistently interrelated by Eqs. (45c). In particular, Eq. (61) implies that $f_{+}$ is asymptotically a linear combination of $\{{\ln\xi_{*}},1\}$, which belongs to the null space of $\hat{L}_{+}$ and therefore most of the information of $f_{+}$ is lost from $\chi_{\gamma}^{(0)}$. Contrastively, the information of $f_{-}$ in Eq. (62) is transmitted by $\hat{L}_{-}$ to $\chi_{\gamma}^{(4)}$ properly, in the sense that both $f_{-}$ in Eq. (62) and $\chi_{\gamma}^{(4)}$ in Eq. (65) contain the same factor $(1-{\mu_{\text{r}}})/{\mu_{\text{r}}}$.
@@ VI.2.2 Second type: uncaged behavior for shorter waves
The integrals needed for evaluation of the DC tensor are still given by Eq. (68) with $m=2$ and $m=4$. For $C_{a}^{\Delta}$ now we use Eq. (76), which implies oscillatory divergence of the integrand for $k\to\infty$. Handling the divergence with the standard technique of convergence factor we find that $I_{a}^{(0)}(\xi_{*})$ vanishes for $\xi_{*}>0$ (the integral reduces to the delta function of $\xi_{*}$). For $m=2$, using the convergence factor $e^{-\epsilon{k}}$ and taking the limit of $\epsilon\to+0$, we obtain
@@ VII Concluding remarks
We have investigated three questions about space–time correlations in a model colloidal liquid. First, we asked whether the displacements have nontrivial correlations even in liquids that are only slightly glassy. The answer is affirmative: this is demonstrated by computing $\mathsf{X}$, the DC tensor defined in Eq. (5), and $\chi_{\gamma}$, the SC defined by Eq. (12), from simulation data of a two-dimensional model colloidal liquid with $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$. The time dependence of these correlations is not linear with regard to the time interval $t_{\Delta}$, as would be expected if the detected correlations were explained as accumulation of many events, but rather described in terms of a similarity variable $\vartheta$ in Eq. (50), indicating the presence of the diffusive correlation length.
Secondly, we asked whether DC and SC are equivalent. The answer is negative: to answer this question, we have derived Eqs. (45c) as a relation between the two kinds of correlations, by treating the displacement field with the label variable formulation. The computed values of DC and SC are then shown to be consistent with Eqs. (45c).
The relation in Eqs. (45c) takes the form of a linear mapping from the DC to the SC, which is expressed by means of two linear operators, $\hat{L}_{\pm}$, of Euler–Cauchy type. This mapping is non-invertible: we can discover $\chi_{\gamma}$ from the components of the DC, $f_{\pm}=X_{\parallel}\pm X_{\perp}$, but cannot recover $f_{\pm}$ uniquely from $\chi_{\gamma}$. In particular, if $f_{+}$ falls into the null space of $\hat{L}_{+}$, the information of $f_{+}$ is lost due to the mapping. In this sense, the DC and the SC are not equivalent.
@@
Interpretation of this information is our third question. It is manifestation of the cage effect at spatiotemporal scales other than usually noticed. The logarithmic behavior of the DC can be explained by assuming a kind of “caged” dynamics for $C_{a}^{\Delta}$, such that the cage effect survives for times longer than $\tau_{\alpha}$ at lengthscales greater than $\ell_{0}$, so that temporal growth of $C_{a}^{\Delta}$ is saturated as is prescribed in Eq. (67). A simple version of such dynamics is exemplified by the elastic modeling in Eq. (52). The dynamics assumed behind Eq. (67), including the elastic modeling, can account for the logarithmic behavior of $f_{+}$ and the non-divergent behavior of $f_{-}$ at the shorter lengthscales. The elastic modeling, however, turns out to be too simple to reproduce the $\phi_{\text{area}}$-dependence of the coefficient $C_{+}$ of the logarithmic behavior in Eq. (66): while the numerical data suggests that $C_{+}$ is an increasing function of $\phi_{\text{area}}$, the elastic modeling makes the opposite prediction, as was noted toward the end of Subsec. VI.3.
@@
Let us now remark on what could be done in the near future within the 2D setup. Since $C_{+}$ seems to carry information on some long-lived aspect of the cage effect at lengthscales greater than $\ell_{0}$, it will be desirable to develop more sophisticated modeling for the dynamics of $C_{a}^{\Delta}$ at such scales. Prediction of $C_{+}$ as a function of $\phi_{\text{area}}$ will serve as a touchstone of such modeling. Viscoelastic modelings of mesoscopic deformations by combination of Maxwell-like phenomenology and the mode-coupling theory (MCT) would be a hopeful direction. While the recent work by Maier et al. in this direction is based on Mori–Zwanzig projection formalism of momentum balance [41], we have been developing a different version of MCT for ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$ on the basis of the (overdamped) Dean–Kawasaki equation [42, 20]. Details on this version of MCT will be reported elsewhere.

### v2

@@ I Introduction
The formulations mentioned above are focused on the mobility of particles and not on their direction of motion. The importance of the directional aspect of motion was noticed by Doliwa and Heuer [11], who demonstrated with Monte Carlo simulations that the displacement correlations in the longitudinal direction behave quite differently from those in the transverse direction. Essentially the same spatial pattern of displacement correlations was also observed independently by Cui et al. [12] in experiments with aqueous suspension of silica spheres. Later, it was noticed that the transverse displacement correlations contain information about glass elasticity [4, 13, 14]. Thus we are interested in tensorial correlations of displacements in glassy liquids, which provides a key topic of the present work.
@@
Motivated by the above considerations, here we concretize our interest in displacement-based statistical quantities by discussing two-particle displacement correlations (DC) and shear strain correlations (SC) in two-dimensional liquids. As is exemplified by the aforementioned experiment by Cui et al. [12], DC and SC can be measured experimentally. Measurement of DC between tracer particles in viscoelastic media has been applied to various soft materials by the name of “two-point micro rheology” [21, 22]. DCs of directly interacting colloidal particles in a suspension with finite density have also been reported [23]. Besides, Illing et al. [20] conducted video microscopy experiments on SC in a glassy liquid. To connect these experimentally (and computationally) measurable statistical quantities with the idea of solid-like liquid theory, now we raise three questions about DC and SC.
Our first question is whether and to what extent these correlations, namely DC and SC, are detectable in liquids that are only slightly dense. To demonstrate the presence of these correlations in liquids is to legitimate the solid-based approach to the liquid dynamics. We focus on such density that the lifetime of “bonds” between the nearest neighbors is barely non-zero and not very long. Although the liquid in such cases does not exhibits macroscopic elasticity at all, we will show numerically an appreciable presence of DC and SC at lengthscales greater than the particle diameter. The time dependence of these correlations will also be investigated.
To make better use of DC and SC as experimentally measurable quantities, we ask further questions about their nature. The second question concerns the possible equivalence between the DC and the SC, as both of them are calculable from the same data of particle displacements. Does the usage of the common ingredient mean that the two kinds of correlations are equivalent? We will answer this question negatively, showing that the SC can be calculated from the DC but not vice versa. In other words, the DCs are more informative than the SCs, which leads to the third question: what is the nature of the extra information contained in DC? This question will be answered, at least partially, by showing that the DCs behave logarithmically for shorter distances, with the coefficient of the logarithmic term being an increasing function of the density. We interpret the nature of this extra information as manifestation of the cage effect.
@@ II.2 Displacement in liquids
The cage effect suggests the possibility that, in dense liquids, hydrodynamic description may be extended to shorter lengthscales close to the interparticle distance [3, 27]. Hydrodynamics means finding a closed set of dynamical equations for slow variables. The long lifetime of the “bonds” can provide grounds for slow variables in dense liquids, such as the temporally coarse-grained velocity of a particle. We emphasize that the situation is different from that of low-density fluids, in which the hydrodynamic velocity field is introduced by spatial coarse-graining of momentum as a conserved quantity. In deriving the Navier–Stokes equation for Newtonian fluids, any confusion between the velocity of a fluid element and the molecular velocity should be severely criticized. In dense liquids, however, the cage effect makes it possible—probably as a crude approximation—to regard the velocity of a particle as the velocity of a “fluid element” containing the particle, with the understanding that fluctuations at the collisional frequency should be filtered out mostly by time-averaging.
@@ II.3 Displacement correlations
Typical behavior of the DCs as functions of $\tilde{\mathbf{d}}$, with $t_{\Delta}$ fixed, is depicted in Fig. 1. Details of the particle model used here [24] will be specified later in Sec. IV. It is seen in Fig. 1(a) that the two components of DC behave quite differently: $X_{\parallel}$ is positive everywhere, while $X_{\perp}$ becomes negative for large $\tilde{{d}}$. This difference was noted in a pioneering work by Doliwa and Heuer [11], many years before the researchers started to notice that the transverse DC may contain information about shear modulus of glassy systems [14, 13]. Essentially the same pattern of displacement correlations is reported, seemingly independently, as an experimental result by Cui et al. [12], who pointed out the presence of “antidrag” region (where $X_{\perp}<0$ in our notation). The directional behavior of DCs can be illustrated pictorially as a pair of vortices [24, 11], as is exemplified in Fig. 1(b).
@@ II.4 Shear strain correlations
The four-petaled SC pattern is often interpreted as a sign of elasticity, showing the response of an elastic media to localized plastic events. Although the observations of the SC pattern in sheared glasses [18, 19] seems to be consistent with this interpretation, it is curious that non-vanishing SC is reported experimentally also in quiescent glassy liquids, being visible even for timescales longer than $\tau_{\alpha}$ [20].
@@ III Relations involving shear strain correlations
In order to answer this question, let us develop a theoretical framework to derive an analytical relation between the DC and the SC. Readers who are more interested in the relation itself than its derivation procedure may skim the theoretical development in Subsec. III.1 and, after checking some basic notation in and around Eqs. (22)–(25), (36) and (41), jump to Eq. (44).
@@ III.2.3 Relation between DC and SC
$=-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{1^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{y}}(\bm{\xi}^{\prime})}\right\rangle-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{2^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{x}}(\bm{\xi}^{\prime})}\right\rangle$ (38)
@@ III.3 Possible nonequivalence between DC and SC
With the relation in Eqs. (45) we can calculate the SC from the DC, but this does not necessarily mean that SC and DC are equivalent. The problem is whether the relation is invertible, making it possible to obtain the DC from the SC. Intuitively speaking, since differentiation is involved in the operators $\hat{L}_{\pm}$, some information in DC is likely to be missing from SC.
The invertibility of the linear operators $\hat{L}_{\pm}$ in Eq. (45) can be discussed by checking their null spaces. The null space of $\hat{L}_{+}$, also known as the kernel of the linear operator [39], is the set of all the solutions $h_{+}$ to the homogeneous equation
@@
with the domain of the function taken as some physically appropriate range of $\xi_{*}$. Noticing that $\hat{L}_{+}$ is essentially the radial part of the Laplacian operator in the $\bm{\xi}_{*}$-space, we can readily find $h_{+}=A+B\log\xi_{*}$ (with $A$ and $B$ denoting arbitrary constants); in other words, the null space is spanned by $\{1,\log\xi_{*}\}$. If $f_{+}\sim h_{+}$, it is mapped to zero by $\hat{L}_{+}$ in Eqs. (45) and therefore the information of $f_{+}$ is lost from $\chi_{\gamma}^{(0)}$. This makes the DC and the SC nonequivalent.
@@
Thus we have derived analytical relations involving the SCs, which allows us to discuss the possibility that the DCs and the SCs are not equivalent due to the null space of $\hat{L}_{\pm}$. Now let us proceed to numerical study of these analytical results, beginning with specification of the particle system in the next section.
@@ IV Specification of the particle system
As the interaction potential $V_{jk}$ between the $j$-th and $k$-th particles (separated by the relative position vector $\mathbf{r}_{jk}$), we adopt the harmonic repulsive potential,
$V_{jk}=\begin{cases}V_{\text{max}}\left(1-\dfrac{\left|{\mathbf{r}_{jk}}\right|}{\sigma}\right)^{2}&(\left|{\mathbf{r}_{jk}}\right|<\sigma)\cr 0&(\text{otherwise})\end{cases}$
@@
Note that here we have adopted a system of monodisperse particles, which makes it simpler to specify the system. This is allowed as we are interested in the validity of the solid-based approach to only slightly glassy liquids.
@@ V Numerical results
We performed numerical simulation of the system of Brownian disks specified in Sec. IV. The system contains $N=4000$ particles, and $L$ is adjusted so as to give three values of the area fraction: $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$. (For higher densities, polydisperse particles would be needed to avoid crystallization, which is out of the scope of the present work.) To mimic the hardcore interaction, we chose $V_{\text{max}}=50$ for $\phi_{\text{area}}=0.50$, $V_{\text{max}}=500$ for $\phi_{\text{area}}=0.60$, and $V_{\text{max}}=5000$ for $\phi_{\text{area}}=0.70$ in units of ${k_{\text{B}}}T$; the time step for numerical integration, $\Delta{t}$, is chosen so as to satisfy $(\Delta{t}/\tau_{0})({k_{\text{B}}}T/V_{\text{max}})=0.1$ where $\tau_{0}=\sigma^{2}/D$.
@@ VI.1 Elastic modeling
In parallel to the term with $D^{\text{c}}_{*}$ in Eq. (52a), the presence of the term with ${\mu_{\text{r}}}D^{\text{c}}_{*}$ in Eq. (52b) means introducing the shear modulus ${\mu_{\text{r}}}{k_{\text{B}}}T/S$ into the model. While this is a natural modeling of DCs in the case of glass solids with plateau modulus or with idealization of $\tau_{\alpha}\to\infty$ [10, 4, 13], it may seem questionable to what extent the elastic modeling is applicable to the case of glassy liquids with modest $\tau_{\alpha}$. Here we regard the elastic modeling as a convenient starting point for searching more sophisticated description of $C_{a}^{\Delta}$ with a wider range of applicability, involving solid-like elasticity at some scale. Recent experiments on confined liquid glyserol [43, 44] and polypropylene glycol [44] may support relevance of such solid-like elasticity in liquids far away from glass transition.
Given the numerical data of the DCs, we focus our attention to the ratio of the coefficients $S$ and $S/{\mu_{\text{r}}}$ in Eqs. (51). In principle, ${\mu_{\text{r}}}$ can be computed from the data in the Fourier space as the ratio of the saturation values of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$:
@@
$X_{\parallel}={\frac{1}{2}}\left({f_{+}}+{f_{-}}\right),\quad X_{\perp}={\frac{1}{2}}\left({f_{+}}-{f_{-}}\right)$ (56)
@@
The analytical expressions of $f_{\pm}$ in Eqs. (61) and (62) allow us to extract information of elasticity from simulational and experimental data of the particle system, on the assumption that the elastic model is quantitatively valid. They may provide useful alternatives to Eq. (53), as $\mathsf{X}(\tilde{\mathbf{d}},t_{\Delta})$ is easier to compute than its Fourier counterpart used in Eq. (53).
@@
It is easy to verify that the above expressions for DCs and SCs are consistently interrelated by Eqs. (45). In particular, Eq. (61) implies that $f_{+}$ is asymptotically a linear combination of $\{{\ln\xi_{*}},1\}$, which belongs to the null space of $\hat{L}_{+}$ and therefore most of the information of $f_{+}$ is lost from $\chi_{\gamma}^{(0)}$. Contrastively, the information of $f_{-}$ in Eq. (62) is transmitted by $\hat{L}_{-}$ to $\chi_{\gamma}^{(4)}$ properly, in the sense that both $f_{-}$ in Eq. (62) and $\chi_{\gamma}^{(4)}$ in Eq. (65) contain the same factor $(1-{\mu_{\text{r}}})/{\mu_{\text{r}}}$.
@@ VI.2.2 Second type: uncaged behavior for shorter waves
The integrals needed for evaluation of the DC tensor are still given by Eq. (68) with $m=2$ and $m=4$. For $C_{a}^{\Delta}$ now we use Eq. (76), which implies oscillatory divergence of the integrand for $k\to\infty$. Handling the divergence with the standard technique of convergence factor, we find that $I_{a}^{(0)}(\xi_{*})$ vanishes for $\xi_{*}>0$ (the integral reduces to the delta function of $\xi_{*}$). For $m=2$, using the convergence factor $e^{-\epsilon{k}}$ and taking the limit of $\epsilon\to+0$, we obtain
@@ VII Concluding remarks
We have investigated three questions about space–time correlations in a model colloidal liquid. First, we asked whether the displacements have nontrivial correlations even in liquids that are only slightly glassy. The answer is affirmative: this is demonstrated by computing $\mathsf{X}$, the DC tensor defined in Eq. (5), and $\chi_{\gamma}$, the SC defined by Eq. (12), from simulation data of a two-dimensional model colloidal liquid with $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$. This is consistent with the experimental observation by Illing et al. [20]. The time dependence of these correlations is not linear with regard to the time interval $t_{\Delta}$, as would be expected if the detected correlations were explained as accumulation of many events, but rather described in terms of a similarity variable $\vartheta$ in Eq. (50), indicating the presence of the diffusive correlation length.
Secondly, we asked whether DC and SC are equivalent. The answer is negative: to answer this question, we have derived Eqs. (45) as a relation between the two kinds of correlations, by treating the displacement field with the label variable formulation. The computed values of DC and SC are then shown to be consistent with Eqs. (45).
The relation in Eqs. (45) takes the form of a linear mapping from the DC to the SC, which is expressed by means of two linear operators, $\hat{L}_{\pm}$, of Euler–Cauchy type. This mapping is non-invertible: we can discover $\chi_{\gamma}$ from the components of the DC, $f_{\pm}=X_{\parallel}\pm X_{\perp}$, but cannot recover $f_{\pm}$ uniquely from $\chi_{\gamma}$. In particular, if $f_{+}$ falls into the null space of $\hat{L}_{+}$, the information of $f_{+}$ is lost due to the mapping. In this sense, the DC and the SC are not equivalent.
@@
Interpretation of this information, contained in $C_{+}$ and $\lambda$ obtainable through Eq. (66), is our third question. It is manifestation of the cage effect at spatiotemporal scales other than usually noticed. The logarithmic behavior of the DC can be explained by assuming a kind of “caged” dynamics for $C_{a}^{\Delta}$, such that the cage effect survives for times longer than $\tau_{\alpha}$ at lengthscales greater than $\ell_{0}$, so that temporal growth of $C_{a}^{\Delta}$ is saturated as is prescribed in Eq. (67). A simple version of such dynamics is exemplified by the elastic modeling in Eq. (52). The dynamics assumed behind Eq. (67), including the elastic modeling, can account for the logarithmic behavior of $f_{+}$ and the non-divergent behavior of $f_{-}$ at the shorter lengthscales. Thus the present work contributes to development of solid-based approach to liquid dynamics, allowing us to extract experimentally verifiable predictions about DC from theoretical modeling of $C_{a}^{\Delta}$, such as the elastic modeling. The elastic modeling, however, turns out to be too simple to reproduce the $\phi_{\text{area}}$-dependence of the coefficient $C_{+}$ of the logarithmic behavior in Eq. (66): while the numerical data suggests (at least for $\phi_{\text{area}}\leq 0.7$) that $C_{+}$ is an increasing function of $\phi_{\text{area}}$, the elastic modeling makes the opposite prediction, as was noted toward the end of Subsec. VI.3.
@@
Let us now remark on what could be done in the near future within the 2D setup. First of all, we would like to invite experimentalists to measure the DC, verify Eq. (66) and thereby evaluate $C_{+}$. Since $C_{+}$ seems to carry information on some long-lived aspect of the cage effect at lengthscales greater than $\ell_{0}$, investigation of its dependence on $t_{\Delta}$ and $\phi_{\text{area}}$ by experiments (both real and simulated) will promote understanding of liquid dynamics at such scales. On the side of theoreticians, it will be desirable that they develop more sophisticated modeling for the dynamics of $C_{a}^{\Delta}$, which, with the aid of the formulae in Subsec. III.2, should allow prediction of $C_{+}$ as a function of $\phi_{\text{area}}$; its experimental verification will serve as a touchstone of such modeling. Viscoelastic modelings of mesoscopic deformations by combination of Maxwell-like phenomenology and the mode-coupling theory (MCT) would be a hopeful direction. While the recent work by Maier et al. in this direction is based on Mori–Zwanzig projection formalism of momentum balance [46], we have been developing a different version of MCT for ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$ on the basis of the (overdamped) Dean–Kawasaki equation [47, 24]. Details on this version of MCT will be reported elsewhere. Here we note, parenthetically, that MCT for tagged particles in Ref. [23] predicts DC with spatial oscillation unobserved in the experiment, and also that two completely different ideas for improvement on MCT for a tagged particle are discussed in Refs. [48, 30].